OpenStudy (toman):

Sixty four men can complete a piece of work in 40 days. In how many days can 32 men complete the same piece of work?

OpenStudy (toman):

Multiply 40 and 64?

OpenStudy (anonymous):

23

OpenStudy (anonymous):

FALK

OpenStudy (anonymous):

6=40 32=x. solve for x

Mehek (mehek14):

64 / 32 = 2 so 32 men do half the work of 64 40 / 2

OpenStudy (toman):

That's not right.

Mehek (mehek14):

no not divide 40 / 2 multiply it by 2 40*2

OpenStudy (toman):

I thought it was: \[\frac{ 40 \times 64 }{ 32}\]

OpenStudy (toman):

Oh okay.